OpenStudy (anonymous):

find the derivatives d/dx (sqrt(x-1/x+1))

OpenStudy (anonymous):

what a pain this is

OpenStudy (anonymous):

you need the chain rule, the quotient rule, and a raft of algebra at the end to simplify ready?

OpenStudy (anonymous):

yupp

OpenStudy (anonymous):

the derivative of \(\sqrt{f(x)}\) is \[\frac{f'(x)}{2\sqrt{f(x)}}\]

OpenStudy (anonymous):

here \[f(x)=\frac{x+1}{x-1}\]so we need to find \(f'(x)\) using the quotient rule

OpenStudy (anonymous):

This problem is easier done by logarithmic differentiation.

OpenStudy (anonymous):

ack typo \[f(x)=\frac{x-1}{x+1}\]so we need to find \(f'(x)\) using the quotient rule

OpenStudy (anonymous):

the x-1 is on the top in my problem

OpenStudy (anonymous):

ok :) just making sure

OpenStudy (anonymous):

@eliassaab yes that is certainly true, but a) i am willing to wager she has not gotten there yet and b) if you leave the answer in that form it will probably not look like what they want

OpenStudy (anonymous):

If you put y=\left\frac {x-1}{x+1} \right)^{1/2}

OpenStudy (anonymous):

\[ y=\left(\frac {x-1}{x+1} \right)^{1/2} \]

OpenStudy (anonymous):

\[\ln (y) =\frac 1 2 \ln (x-1) - \frac 1 2 \ln (x+1) \]

OpenStudy (anonymous):

in any case i have to run because the twilight zone is coming on, but you need \[f'(x)\] which in this case is not too hard, \[\frac{(x+1)-(x-1)}{(x+1)^2}=\frac{2}{(x+1)^2}\]

OpenStudy (anonymous):

so your final job is to write \[\frac{\frac{2}{(x+1)^2}}{2\sqrt{\frac{x-1}{x+1}}}\] and then have an algebra party

OpenStudy (anonymous):

ok im sorry but im really confused now

OpenStudy (anonymous):

\[ \frac {y'} y =\frac 1{2 (x-1)}-\frac 1{2 (x+1)} \\ y'= y \left ( \frac 1{2 (x-1)}-\frac 1{2 (x+1)} \right) \]

OpenStudy (anonymous):

i am sure @eliassaab way will work swifty land nicely, but the devil will still be in the algebra if you know logarithmic diff, use it

OpenStudy (anonymous):

i dont think i really do so can we start from the top and go through it pplleeassee

OpenStudy (anonymous):

look at the first post i wrote to give the idea of what you have to compute i did the necessary computation what is left is the algebra to make \[\frac{\frac{2}{(x+1)^2}}{2\sqrt{\frac{x-1}{x+1}}}\] look more reasonable hold on i need to log off and log on again back in 2 minutes

OpenStudy (anonymous):

Final simplified answer \[ y'=\frac{1}{2} \sqrt{\frac{x-1}{x+1}} \left(\frac{1}{x-1}-\frac{1}{x+ 1}\right)=\frac{\sqrt{\frac{x-1 }{x+1}}}{x^2-1} \]

OpenStudy (anonymous):

An equivalent answer will be \[y'=\frac{1}{\sqrt{\frac{x-1}{x+1}} (x+1)^2} \]

OpenStudy (anonymous):

if you got this, let me know if you want to start at the top we can do that too

OpenStudy (anonymous):

i think i got it

OpenStudy (anonymous):

\[d/dx \left[ \left( x ^{2+1} \right)\sin \sqrt{x} \right]\]

OpenStudy (anonymous):

2+1?

OpenStudy (anonymous):

supposed to be \[x^{2}+1\]

OpenStudy (anonymous):

does that make sense?